Accessibility Assistance

Skip to Content

Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
August 3, 2010

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Debra Fetzer, Michael Jay, Tom Minton, Trish Ruwell, Evan Wack

Guest: Rachel Schipper
  1. Will and Cynthia will be handling orders for the next two weeks.
  2. Cathy Martyniak will be out of the office for the rest of the week, if you are trying to reach her.
  3. Jack Waters recently had bypass surgery. He is recuperating.
  4. Peer-to-Peer this Friday, August 6, 8:30-12PM at the Reitz Union. Many topics of interest: Green IT, IT Wiki, OCS, Shibboleth, security, etc. Rachel will not be able to make it, but would appreciate a summary of the topics discussed. Will has asked for teleconferencing. Otherwise, some of us will need to go and report back.
  5. Division meeting feedback - Will to discuss at TSS tomorrow.
Ongoing IT Updates

  1. West computer replacement - plan for next week: install 25 computers per day to get it done in a week's time. West staff will be informed of our plans. Work force to be divided up into: runners, installers and data entry. Need 8 folks to install, 4 to run and 2 for data entry. Campus cab will be at Eastside at 8:30AM (hold 6 people). Facilities van will also be there (holds 6). Work to stop at 11:30AM, vans return to Eastside at 12:30PM. Old machines will be surveyed out of Facilities. The West cleaning crew has been vacuuming under the tables. Cynthia will use her cell phone as a contact for Campus Cab. Folks can volunteer for a position; otherwise, Evan will assign you one.
  2. Strange ongoing networking problem - weird issues the last two weeks. Unsure if the issue is with the network, policy or computers. Right now it's looking like the Domain Controller. The problem is only affecting Win7 machines. Evan is quite sure that it is not the image. The networking problem in West this morning was an issue with the DHCP server: CNS forgot to add the West IP ranges. The ongoing network issue affects Win7 on DHCP. Problem did not exist until CNS changed their DHCP server. CNS is currently watching the traffic flow. There is no failure outside of the public OU. As a work-around, we may take all group policy and put it in local policy. The only other option is to stick with XP.
  3. Are all items removed from the East vaults? Not yet, there are still things awaiting survey.
  4. Pharos guy is here and Josh is training with him today on firmware updates. Josh is learning how to troubleshoot. One of the Canon reps stopped by here today. He said the new software should be released by them in two weeks. Lag time has improved. Duplexing is done. Drivers have been added for varying types of USB drives. Laptop printing is working, but has not been rolled out yet. Topic of how to announce it to be addressed at Branch Chairs. Need to include Barbara Hood.
  5. 1A microphones - the AC noise has dramatically decreased. Evan contacted the AV folks and they will be scheduling a time to return and reprogram the ceiling mics.
  6. Public computer image on hold until networking problem is resolved.
  7. Time Limited Workstations - Evan and Josh demoed the time-limited software to Stacey Ewing and Jana Ronan. They are very pleased with it. Josh will install it tomorrow. They have decided to block in one, two, three or four hour blocks. They would like on-the-floor training.
  8. Privilege Authority - has been given to VIVO folks.
  9. Staff migration - on hold.
  10. Diane Bruxvoort (John Ingram's replacement) - what is the status of her computer? She needs to be set up as she will be in town very soon. John Ingram needs to be moved over to Special Collections.
  1. Winston is in Atlanta at ASERL.
  2. GovDocs meeting - they liked the screen shots we showed them. Winston will be showing it at ASERL.
  3. Tom is working on three FCLA reports.
  4. Shruti is working on ITS-related tickets.
  5. Michael Jay - the Jason Fleming macro is done. The Paul Lightcap order tracking macro is on hold - waiting for input from him, but he is in transition right now.
  6. Maureen Kelly - new OLE programmer starts August 27. She will be located in Access Services, 2nd floor West. We need to order a computer and printer for her. Need to get her a phone. It is a two-year, lime-limited position. We will also be hiring a time-limited programmer to work with her in March.
Network/Server Administration
  1. Sharepoint is going along, although Logan has been busy dealing with the networking issue. He is scheduled to meet with Steve Carrico and Raimonda this week to work on their template. Logan will wait for Chris to return from vacation. After that, we can roll it out to other departments that have expressed interest. Rachel notes that Acquisitions is having some staff issues, so you may want to double check. Cataloging has also expressed an interest, so we can always start with them if Acquisitions is not ready.
  2. Virtualization - moving along. The GIS image has been moved across. We will try to upgrade it to 2K8 tomorrow. Need to figure out Hashem's Maps application.
  3. Staff printing - will get to it next week when Amy returns. Evan mentions renaming the print queues - trying establish a naming convention.
  4. App server - on schedule, waiting for Chris Nicolich to return.
  5. VIVO - migration finalized. Server set up for conference on August 12. Logan spoke with Grace regarding the fiscal issue.
  6. MSL - servers offline. Shutting down UFDC server on Friday. This will make available a server for Altiris. Need to figure out what to do with the rest of the boxes.
  1. Going well - Cynthia had four tickets. Three of them are now done. Michael helped her figure out an issue today.
DLC General Issues
  1. OCR issues. Sending the ABBY key to Laurie Taylor tomorrow. Dongle to be installed on DLC machine.
  2. Troubles with OCR automation on OCR4 - SAN permissions issue. Working on it.
  3. UFDC - looking good. No complaints. Just need documentation.
Web Projects
  1. Lots of URL changes.
  2. DLC has its own 4th level domain. Now it doesn't show in searches. It also breaks Google analytics. Rachel notes that Judy has often mentioned the long length of the DLC links. She would like them shortened. It would have to be done on the back end. Will need to discuss this with Mark Sullivan.
  3. August 9 & 10 - lots of URL changes will be implemented. Links and search engines need to be redone.
  4. LibGuides and Course Guides to change to Campus Guides. This removes them from the Library. Isn't the thought to steer them to the Library?